### v1

# Open Vocabulary Semantic Scene Sketch Understanding

Abstract
We study the underexplored but fundamental vision problem of machine understanding of abstract freehand scene sketches. We introduce a sketch encoder that results in semantically-aware feature space, which we evaluate by testing its performance on a semantic sketch segmentation task. To train our model we rely only on the availability of bitmap sketches with their brief captions and do not require any pixel-level annotations. To obtain generalization to a large set of sketches and categories, we build on a vision transformer encoder pretrained with the CLIP model. We freeze the text encoder and perform visual-prompt tuning of the visual encoder branch while introducing a set of critical modifications. Firstly, we augment the classical key-query (k-q) self-attention blocks with value-value (v-v) self-attention blocks. Central to our model is a two-level hierarchical network design that enables efficient semantic disentanglement: The first level ensures holistic scene sketch encoding, and the second level focuses on individual categories. We, then, in the second level of the hierarchy, introduce a cross-attention between textual and visual branches. Our method outperforms zero-shot CLIP pixel accuracy of segmentation results by 37 points, reaching an accuracy of on the FS-COCO sketch dataset. Finally, we conduct a user study that allows us to identify further improvements needed over our method to reconcile machine and human understanding of scene sketches.
1 Introduction
Even a quick sketch can convey rich information about what is relevant in a visual scene: what objects there are and how they are arranged. However, little work has been devoted to the task of machine scene sketch understanding, largely due to a lack of data. Understanding sketches with methods designed for images is challenging because sketches have very different statistics from images – they are sparser and lack detailed color and texture information. Moreover, sketches contain abstraction at multiple levels: the holistic scene level and the object level. Here we explore the promise of two main ideas: (1) the use of language to guide the learning of how to parse scene sketches and (2) a two-level training network design for holistic scene understanding and individual categories recognition.
Freehand sketches can be represented as a sequence or cloud of individual strokes, or as a bitmap image. As one of the first works on scene sketch understanding, we target a general setting where we assume only the availability of bitmap representations.
We also aim at the method that can generalize to a large number of scenes and object categories. To this end, we build our sketch encoder on a Visual Transformer (ViT) encoder pre-trained with a popular CLIP [43] foundation model (Fig. 1). We propose a two-level hierarchy in our network, where the two levels (“Holistic” and “Category-level”) share the weights of our visual encoder. The first level focuses on ensuring that our model can capture holistic scene understanding (Fig. 2: I. Holistic), while the second level ensures that the encoder can efficiently encode and distinguish individual categories (Fig. 2: II. Category-level). We avoid reliance on tedious user per-pixel annotations by leveraging sketch-caption pairs from the FS-COCO dataset [8], and aligning the visual tokens of sketch patches with textual tokens from the sketch captions, using triplet loss training. We strengthen the alignment by introducing sketch-text cross-attention in the second level of the network’s hierarchy (Fig. 2: g.). Additionally, we introduce a modified self-attention computation to the visual transformer encoder used in both layers, inspired by recent work by Li et al. [32].
We conduct a comprehensive evaluation of our method comparing it with recent language-supervised image segmentation methods [58, 43, 32], fine-tuned on the FS-COCO dataset. We show that our approach outperforms with a large margin all existing methods on the task of freehand sketch segmentation. We also compare with a previous fully supervised work on scene sketch segmentation [19], trained on a semi-synthetic set of sketches composed of individual category sketches. We demonstrate that their work does not generalize well to freehand scene sketches from the FS-COCO dataset. Our method demonstrates consistent performance and similarly outperforms [19] on a dataset of freehand sketches provided by Ge et al. [19].
Finally, our analysis reveals that although our model consistently produces robust segmentation results across the majority of sketches, there are a few challenging sketching scenarios for our method. We select a subset of representative sketches for each scenario and collect multi-user annotations. We then carefully assess our approach by comparing its performance with that of human participants, drawing insights to guide future work.
In summary, our contributions include: (1) a two-level network design used during training, focusing on holistic scene sketch understanding and category disentanglement, (2) the first language-supervised scene sketch segmentation method, (3) per pixel segmentation annotations of 975 sketches from the FS-COCO dataset, and (4) multi-user annotations of a subset of distinct groups of sketches.
2 Related Work
2.1 Unsupervised and weekly supervised image semantic segmentation
The need for pixel-wise segmentation limits the amount of data that fully supervised segmentation models [35, 64, 5, 6, 1, 17] can use for training, as such annotations are costly to collect. This in its turn limits the generalization properties of models trained with pixel-level annotations. To avoid the need for extensive annotations, unsupervised [62, 7, 23, 25, 38], semi-supervised [39, 72] and weakly supervised [40, 56, 58, 37, 13, 12, 36, 69] methods were proposed.
Our method belongs to the group of weakly supervised methods based on text annotations only [58, 37, 13, 12, 36, 4], such methods are not limited to a fixed set of categories and therefore are referred to as open vocabulary semantic segmentation methods. Image methods typically rely on the spatial proximity of semantically similar pixels. This is less applicable in the sparse and largely monochromatic landscape of freehand sketches. For example, recent GroupViT [58] and SegCLIP [37] use learnable group tokens and semantic group modules to aggregate low-layer pixel features. In our work, we propose a two-level training architecture taking sketch specifics into account.
2.2 Sketch semantic segmentation
The majority of works on semantic sketch segmentation focus on single-category sketches. Some of these works treat sketch as a bitmap image [70, 31, 71], but most leverage stroke-level information directly [21, 27, 57, 42, 11, 48, 22, 55, 60, 41, 66] or as a segmentation refinement step [31, 71]. All these works are fully supervised except for [41], which segments sketches of a given category provided at least one segmented reference sketch.
Semantic scene sketch segmentation [51], and more broadly scene sketch understanding, is underexplored to a large extent due to lack of data. The lack of data is typically addressed by introducing semi-synthetic sketch datasets. The SketchyScene dataset [73] consists of 7,264 sketch-image pairs, obtained by arranging clip-art sketches in alignment with a reference image. SketchyCOCO dataset [18] is generated from COCO-Stuff [3] by semi-automatically arranging freehand sketches of individual categories. Ge et al. [19] introduced their own semi-synthetic scene sketch dataset and adopted a DeepLab-v2 [5] architecture to the scene sketch segmentation task. SketchSeger [59] proposed an encoder-decoder model based on hierarchical Transformers, trained with a stroke-based cross-entropy loss on semi-synthetic scene sketches formed by combining sketches from the QuickDraw dataset [21]. Zhang et al. [63] proposed an RNN-GCN-based architecture trained on annotated freehand scene sketches. However, neither the dataset nor the code have been released. We do not require stroke-level information or pixel-wise segmentation of the training data, and leverage the FS-COCO dataset [8] of freehand sketches with their textual descriptions.
2.3 ViT-CLIP and sketch
We build our encoder on a ViT (Vision Transformer) encoder pre-trained with CLIP (Contrastive Language-Image Pre-training) [43]. CLIP is a model trained on roughly 400 million image-text pairs to embed images and text in a shared space. It uses ViT as a visual branch (image) encoder. A ViT encoder pre-trained with CLIP (ViT-CLIP) is used in a range of sketch-related tasks: sketch and drawing generation [49, 16, 54, 53], 2D image retrieval [47, 8, 45], object detection [9], 3D shape retrieval [50, 29, 61, 30, 2], 3D shape generation [65].
While some works use ViT-CLIP purely pre-trained on images, many fine-tune the encoder on sketches for downstream tasks. Some works fine-tune all weights of the encoder [47, 2], some fine-tune Layer Normalization layers only [8], and some rely on prompt-learning [26, 68] or the combination of the latter two [45, 9]. In our work, we also rely on fine-tuning with visual prompt learning and Layer Normalization layers updates. Unlike previous methods targeting sketch inputs, we additionally leverage a two-path ViT architecture, inspired by Li et al. [32].
3 Method
As we mention in the introduction, we build a sketch encoder such that the semantic meaning of individual stroke pixels can be inferred from its feature embeddings. Building on the ViT encoder, pre-trained CLIP [43] model, we fine-tune a modified encoder architecture with a network consisting of two levels: Holistic scene understanding and individual category recognition. We start by describing the first level of our network (Fig. 2 I.) and introducing the architecture of our visual encoder (Fig. 2 c.). We then describe our strategy to improve the model’s ability to understand individual categories (Fig. 2 II.).
3.1 Holistic scene sketch understanding
The architecture in the first level (Fig. 2: I. Holistic) is similar to the architecture of the CLIP model [32]. We freeze the weights of the textual encoder and fine-tune the modified architecture of the vision encoder (Sec. 3.1.1). The CLIP model is trained with a contrastive loss, ensuring that the embedding of images and corresponding captions are closer in space than embeddings of images and captions of other images. While our training has a similar goal, we train with a triplet loss with hard triplet mining, as we found it to be more beneficial with the batch size we use:
Here, a holistic visual scene sketch embedding VST (Visual Scene Token) serves as an anchor. An encoding of the matching sketch caption (Caption Scene Token) serves as a positive sample, and an encoding of the most dissimilar scene caption serves as a negative sample . We set the margin to a commonly used value of . The number of triplets is equal to the number of samples in a batch.
3.1.1 Visual encoder
The input scene sketch is divided into non-overlapping patches, which are flattened and linearly projected into the feature space. Concatenating with positional encodings, we obtain one token per patch. Additionally, we add a set of learnable tokens, , referred to as visual prompts [10]. Finally, these tokens are also augmented with a special token that encodes holistic sketch meaning, VST (Visual Scene Token). Note that in the context of classification, a CLS token has a similar role to our VST token. Therefore, the input to the vision encoder is , where .
Attention computation in the visual encoder
It was observed by Li et al. [32] that CLIP-predicted similarity maps between image and text features emphasize background regions rather than areas that correspond to a category in the text embedding.
To address this issue, they proposed an alternative configuration of the visual transformer that does not require training or fine-tuning the original model. They demonstrated that this configuration leads to improved performance in open vocabulary segmentation tasks.
We performed a similar experiment with CLIP features for sketch inputs: The similarity maps in the second row of Fig. 3 show the poor ability of CLIP features to identify target categories.
Therefore, we follow Li et al. [32] and use their two-path configuration of the vision transformer. However, we use it not only for inference, but also incorporate this two-path configuration directly into our network training, as we find it more beneficial. We provide a detailed analysis in Sec. 4.5.1.
The first path represents the original vision encoder where identical blocks are repeated times. Each block consists of Layer Normalization (LN), followed by Multi-Head Self Attention (MHSA), another LN and Fead Forward Network (FFD).
The second path blocks contain a modified attention computation in MHSA, dubbed as v-v self-attention, where Keys and Queries are ignored, and self-attention is computed using only Values, :
In addition, blocks in the second path do not include the second LN and FFN layers. Finally, in the second path, the input to the v-v multi-head attention is always the features from the original path. We use the output from the second path during training and inference.
As shown in Fig. 3 third row, the v-v attention results in feature representations that accurately represent distinct semantic entities present in the scene sketch.
3.2 Categories disentanglement
Given the sketch caption we automatically identify individual categories and generate a set of textual prompts of the form “A sketch of *” (Fig. 2b.). Each of these textual category prompts is encoded with the CLIP text encoder into (Caption Category Token).
We then compute the per-patch cosine similarity between the class embeddings and the scene sketch patch embeddings , defined as:
where is the patch index and is an index of a category (e.g. tree). The resulting similarity matrix represents the category label probabilities for each individual patch (Fig. 2d.). To generate a pixel-level similarity map, we reshape each and then upscale to the dimensions of the original scene sketch using bi-cubic interpolation [52]. By multiplying these per category maps with the input scene sketch, as shown in Fig. 2e., we obtain disentanglement into individual sketch categories.
3.2.1 Thresholding with a learnable parameter
Only pixels with similarity scores above a certain threshold are retained at this step (Fig. 2f.). We make the threshold learnable, eliminating the need for manual tuning. More importantly, the threshold value increases over epochs as the model becomes more confident in its predictions, allowing the model to obtain strong disentanglement performance.
3.2.2 Visual Encoder with Cross Attention
The features of individual category sketches are extracted with the visual encoder identical to the one used in the holistic scene sketch level understanding of our network, described in Sec. 3.1.1, up to one difference.
We enhance the interplay between textual and visual domains through the introduction of cross-attention. Namely, in 7th, 10th, and 12th layers in the MHSA, we feed CCT token from the textual encoder representing a target category to the linear projection for the queries. This enables the model to leverage category token embedding from the textual domain to update the sketch token embedding. This results in a better text-to-sketch alignment for individual categories and subsequently improves sketch semantic segmentation. Our ablation study in Tab. 4 underscores the efficacy of this cross-attention strategy.
3.2.3 Text-sketch category-level alignment
We train with a triplet loss, , so that the category sketch embedding, VCT (Vision Category Token), is used as an anchor, the matching embedding of the category prompt is used as a positive sample and the embedding of the prompt of the most dissimilar category is used as negative. We use the VCT from multiple encoder layers: .
3.3 Efficient CLIP fine-tuning
The two levels (holistic and category) are trained jointly, using the total loss
We leverage the generalization properties of the pre-trained foundation model through careful fine-tuning. We freeze all the weights apart from weights of LN, as was proposed in [15], and we use learnable visual prompts, as was proposed in [26]. We introduced visual prompts in Sec. 3.1.1. We also train linear layers which take part in cross-attention computation.
3.4 Inference
Our network design allows segmentation for different sets of categories. Given a desirable set of categories for a given sketch, we obtain sketch segmentation by applying all the steps of our network up to the calculation of pixel-category similarities (Fig. 2e.), followed by upscaling of similarity maps for each category, as discussed in Sec. 3.2. To assign segmentation results we assign to each pixel a label that yields the highest similarity value across category similarity maps , where is an index of a category.
If we want to isolate just a few categories in the sketch, we can use the thresholding strategy that we use during training to isolate the pixels of a given category (Fig. 2f.). We used this strategy to obtain visualizations in Fig. 1, with a threshold value of that we found to be optimal on the test set of sketches. We do not use the learned value from the training, as during training the model does not have to select all the pixels of the given category, but only those that are sufficient to confidently predict the category label. We provide in-depth discussion in the supplemental.
4 Experiments
4.1 Training and test data
For training and testing, we use the sketch-caption pairs from the FS-COCO [8] dataset. The dataset comprises 10,000 sketch-caption pairs, associated with reference images from the MS-COCO [33] dataset. The sketches are drawn from memory by 100 non-expert participants. The reference image was shown for 60 seconds, followed by a 3-minute sketching window.
Training/Validation/Test splits We first selected 500 sketches with distinct styles from five participants. We then randomly sample 5 sketches from each of the remaining 95 participants for validation (a total of 475 sketches). We use the remaining 9025 sketches for training.
Test/Validation set annotations One of the co-authors manually annotated test and validation sketches, relying on reference images and category labels from the MS-COCO [33] dataset. We assign each stroke a unique category label. Candidate category labels are extracted from MS-COCO image captions rather than sketch captions to obtain richer ‘ground-truth’ annotations. Our test set contains 185 different object classes, with an average of 3.54 objects per sketch.
4.2 Evaluation metrics
We use standard metrics, commonly used in sketch segmentation literature [24, 57, 63].
Mean Intersection over Union ():
evaluates the average of the ratios between the intersection and the union of ground truth and predicted labels over all categories.
Pixel Accuracy (): measures the ratio of correctly labeled pixels to the total pixel count in a sketch.
Stroke Accuracy (): evaluates the percentage of correctly classified strokes to total strokes per sketch. A stroke label is determined by its most frequent pixel label.
4.3 Implementation Details
We implemented our method in PyTorch and trained on two GB Nvidia RTX A GPUs. We built on CLIP [43] with a ViT backbone using ViT-B/16 weights. The input sketch image size is set as . We use 3 learnable visual prompts. We use AdamW optimizer with a learning rate of , and train the model for epochs with a batch size of . We pick a checkpoint based on the mIoU performance on the validation set. We provide more discussion on the checkpoint choice in the supplemental.
4.4 Comparison against state-of-the-art
4.4.1 Comparison with methods relying on the availability of pixel-level annotations
We first compare with several recent methods for image segmentation that similarly to us utilize either CLIP as a backbone: DenseCLIP [44] and ZegCLIP [69], or more recent foundational backbones Grounding-DINO [34] and SAM [28], used in Grounded-SAM [20]. These methods require a certain amount of pixel-level annotated examples. We also compare to a recent fully supervised method [19] for scene sketch semantic segmentation, which is trained on a dataset of semi-synthetic sketches. Such sketches are obtained as a superposition of freehand category-level sketches. We refer to this method as LDP (Local Detail Perception).
Tab. 1 shows that neither of the considered methods generalizes well to freehand scene sketches with a large number of categories.
| Methods | |||
| ZegCLIP [69] | |||
| DenseCLIP [67] | |||
| Grounded-SAM [20] | |||
| LDP [19] | |||
| Ours | 73.48 | 85.54 | 87.02 |
4.4.2 Comparison with language-supervised methods
Next, we compare with several recent methods targeting semantic segmentation with ViT encoders and image-text supervision: GroupViT [58] and SegCLIP [37]. Additionally, we compare with CLIP [43], as well as CLIP Surgery [32] that introduced the usage of v-v-attention at inference time.
Zero-shot In Tab. 2, we first compare the performance of our method with the zero-shot performance of these methods. It shows that image segmentation methods do not generalize well to freehand sketches.
Fine-tuning We fine-tune each of the methods on our training set, by updating all their weights. Since such fine-tuning might be sensitive to a learning-rate choice, we perform several runs with several settings of learning rate parameters. We chose the setting for each method that results in the best performance on our validation set. The fine-tuned methods are marked with stars.
Table 2 shows that our method outperforms all considered baselines, and surpasses the best-performing baseline CLIP Surgery by a substantial margin of , and points in score, and , respectively. In Sec. 4.5.1, we evaluate various elements of our architecture and their contribution to overall performance.
| Methods | |||
| CLIP [43] | |||
| GroupViT [58] | |||
| SegCLIP [37] | |||
| CLIP_Surgery [32] | |||
| CLIP | |||
| GroupViT | |||
| SegCLIP | |||
| CLIP_Surgery | |||
| CLIP_Surgery | |||
| Ours | 73.48 | 85.54 | 87.02 |
Fig. 5 shows the qualitative comparison between our method and the CLIP Surgery. We provide additional visual comparisons in the supplemental.
4.4.3 Generalization ability of our method
Here, we evaluate our method on an additional dataset of 50 freehand sketches provided and annotated by Ge et al. [19]. Tab. 3 shows that our model again demonstrates superior performance on this dataset over the method [19], fully supervised on semi-synthetic sketches. We do not compute Acc@S as sketches are only available as bitmap images. This experiment highlights that short language captions can be efficiently used for training, eliminating the need for expensive and time-consuming per-pixel annotations.
The lower mIoU values on these sketches than on FS-COCO sketches can be explained by (1) on larger average number of categories in them ( categories per sketch) than in our FS-COCO test set ( categories per sketch); (2) domain gap. The sketches from [19] contain symbolic representations of objects (see inset
4.5 Ablation study
4.5.1 Importance of individual components
We perform an ablation analysis to assess the importance of each component in our architecture. Tab. 4 shows the performance of the complete model with individual elements removed. We discuss them in order of impact on overall performance.
v-v attention First, we show the importance of the v-v attention, by substituting our dual path v-v attention-based ViT encoder with the original configuration used in the CLIP model (w/o v-v attention).
Two-level network architecture We keep only the first level of holistic scene understanding of the network (Fig. 2 I.). This architecture is similar to CLIP Surgery, but is supervised with the triplet loss and is fine-tuned using learnable visual prompts and updates only LN layers. Tab. 4 (w/o category-level) confirms that two-level network architecture, along v-v attention, is central to the superiority of our model.
Thresholding We perform an experiment where instead of thresholding we weight each pixel according to cosine similarity scores in maps (Tab. 4 (w/o thresholding)). The learnable threshold more efficiently filters out irrelevant pixels, forcing the model to learn superior disentanglement of individual categories.
Holistic scene encoding Removing the global loss, given by Eq. 1, similarly results in the performance drop (w/o Global Loss). This shows the mutual importance of the two levels of our network.
Cross-Attention Cross attention also substantially contributes to performance. If we use a ViT encoder at the second level of the network (category level), identical to the one used at the first level (holistic level) (Fig. 2c.), then the performance drops by a noticeable points in the mIoU score (Tab. 4 (w/o cross-attention)).
Multi-layer features in the triplet loss Tab. 4 (w/o cross-attention) shows that using features from multiple layers () in the category-level triplet loss is beneficial over using only the features from the last layer ().
| Model | |||
| w/o v-v attention | |||
| w/o category-level | |||
| w/o thresholding | |||
| w/o global Loss | |||
| w/o cross-attention | |||
| w/o multi-layer Loss | |||
| Ours-full | 73.48 | 85.54 | 87.02 |
4.5.2 Efficient fine-tuning
Fig. 6 shows the comparison of different fine-tuning strategies. We obtain the best results by combining fine-tuning of LN (Layer Normalization) layers and the addition of learnable tokens. Adding more or less tokens degrades the performance Fig. 6b.
5 Human-model alignment
Fig. 7 shows that for the majority of sketches in our test set from the FS-COCO dataset, our model correctly labels more than pixels.
In this section, we investigate (1) which sketches are likely to get low segmentation accuracy and (2) how the prediction of our model compares with human observers across different groups of sketches.
5.1 Sketch groups
We identified four distinct sketch groups that are challenging for our model: (1) Ambiguous sketches: sketches where it might be hard even for a human observer to understand an input sketch; (2) Interchangeable categories: sketches containing multiple objects with labels that can interchange each other, like ‘tower’ and ‘building’, or ‘girl’ and ‘man’; (3) Correlated categories: sketches with categories that typically co-occur in scenes, e.g., ‘train’-‘railway’ and ‘airplane’-‘runway’; and (4) Numerous-categories: sketches with six or more categories.
We supplement these four groups with sketches where our model labels correctly more than of pixels: (5) Strong performance.
5.2 User study setting
Data
We sample 5 sketches for each of the first 4 categories and 10 sketches for the 5th category. We visualize selected sketches in the supplemental material.
Participants We recruited 25 participants ( male) for this study. Each participant was randomly assigned six sketches: 1 from each of the first 4 groups and 2 from the 5th group, such that every sketch was annotated by five unique participants.
Study Procedure Participants were presented with one sketch and one object category at a time and were not able to see their previous annotations. Sketch-category pairs were interlaced, to reduce the effect of memorizing their previous annotation on a certain sketch. The annotation interface enabled precise pixel-level segmentation by allowing participants to “paint” over each sketch using a brush with an adjustable radius. Participants could also use the eraser to correct erroneous annotations. Once a participant has moved to a new sketch-category pair, they were not able to change their previous annotations.
5.3 User study analysis and Future work
‘Human’ segmentation
For each sketch, we generate one ‘human’ segmentation using a majority vote. For each pixel and each label, we computed the percentage of annotators that assigned a given label. We then assigned to each pixel the label that was provided most frequently to that pixel by different annotators. In cases where there were multiple labels were provided equally often for a pixel, we randomly sampled one of these labels.
Analysis First, we observed that on sketches that did not fall into any of the challenging categories, our model almost reaches human-level performance, with a negligible gap of points on average (Fig. 8 Strong).
Interestingly, given a label humans can correctly identify sketch pixels even in the presence of ambiguity (Fig. 8 Ambiguous). While none of the models currently match human performance on ambiguous sketches, our model surpasses the other methods by a noticeable margin, demonstrating the effectiveness of our two-level training architecture.
The performance across semantically interchangeable categories is uniform amongst the three language-supervised models. This potentiality can be alleviated by proposing solutions that assign labels jointly.
On sketches with correlated categories our model and ClipSurgery perform similarly, highlighting the inherent limitation of training using language supervision. For a few such categories, one might need to further fine-tune the model relying on sketches of isolated categories.
Our model represents a substantial improvement over current alternatives, surpassing them by more than points. Future work should seek to improve alignment with human sketch understanding, especially on sketches with more than six categories (Fig. 8 Numerous).
6 Conclusion
While focusing on the task of sketch segmentation, we introduced a strategy to train a ViT encoder that results in the feature space with good semantic disentanglement. Such feature spaces contribute towards improving machine understanding of abstract freehand sketches and underpin a range of downstream tasks such as communication and creative pipelines. In light of the latter, it can enable more potent tools for conditional generation and retrieval. In psychology, sketches are used to analyze cognitive functions. This can be facilitated by the availability of robust sketch understanding tools. Importantly, we for the first time demonstrated how language supervision can be used for the task of scene sketch segmentation. Finally, we conducted a comprehensive analysis of our model’s performance, identifying research directions to further align the understanding of sketches by humans and machines.
S1 Overview of the supplementary material
In Sec. S2.1, we provide additional visual comparisons of the results obtained with our method versus results obtained with the state-of-the-art language-supervised image segmentation methods.
In Sec. S3.1, we analyze segmentation accuracy per category.
In Sec. S2.3, we further investigate the generalization properties of our method and how it compares with fully-supervised methods.
In Sec. S4.1, we provide a detailed analysis of the benefit of using cross-attention.
In Sec. S4.2, we analyze different models’ performance depending on the choice of a checkpoint: the last checkpoint versus the checkpoint optimal on the validation set.
In Sec. S4.3, we discuss in detail the choice of a threshold value for segmenting out pixels corresponding to individual categories.
S2 Additional analysis of our method performance
S2.1 Additional qualitative comparisons
In the main paper, we show in Tab. 2 a numerical comparison of the segmentation results obtained with our method and the segmentation results obtained with the state-of-the-art language-supervised image segmentation methods. Also, in the main paper, in Fig. 5, we show a comparison of our model with CLIP_Surgery, where CLIP Surgery represents the fine-tuned CLIP_Surgery [32] model with v-v self-attention introduced at both training and inference stages. Here, in Figs. S9 and S10, we provide an additional visual comparison between our method and state-of-the-art language-supervised image segmentation methods: GroupViT [58], SegCLIP [37], CLIP_Surgery [32], fine-tuned on the FS-COCO dataset. The fine-tuned versions of these models are denoted as GroupViT, SegCLIP, CLIP_Surgery, respectively. In Figs. S9 and S10, we show segmentation results and the error maps (in red), which visualize incorrectly labeled pixels for each method.
S2.2 Segmentation accuracy analysis by category
In this section, we analyze segmentation accuracy per category in both the train and test sets. We show in Fig. S11 the pixel accuracy (Acc@P) for each selected object category. For the figure, we selected categories that appear more than ten times in FS-COCO dataset [8] captions. First, we can see that the segmentation accuracy is smoothly distributed across different categories.
Next, we investigate whether more frequent categories are more likely to be labeled accurately. To evaluate this, we approximate the frequency of a category by counting its occurrence in both the train and test sets, then consider only categories that appear in the test set. We plot with a green and red lines in Fig. S11 the train and test sets category frequency, respectively.
The figure clearly shows a lack of correlation between the frequency of category occurrence and its segmentation accuracy.
We further evaluate it numerically by computing the correlation between , the pixel accuracy (Acc@P) of each category, and the occurrence frequency of this category:
where is the number of categories in the test set.
The resulting correlation coefficients for both train and test sets are 0.16 and 0.14, respectively. This suggests a very weak accuracy-frequency correspondence, indicating that our model is not biased toward more frequently occurring categories. We hypothesize that this is in part due to our careful fine-tuning strategy, which prevents over-fitting. Therefore, the model efficiently leverages pre-training on a large image dataset.
S2.3 Additional discussion on synthetic vs. freehand sketches
In the Sec. 4.4.3 in the main paper, to better understand the generalization properties of our model, we evaluated our method trained on the sketches from the FS-COCO dataset [8] on the freehand sketches from [19]. Here, we provide additional analysis of generalization properties.
S2.3.1 Generalization to sketches consisting of clip-art-like object sketches
Here, we additionally evaluate our method on the SketchyScene [73] dataset. The SketchyScene [73] dataset contains 7,264 sketch-image pairs. It is obtained by providing participants with a reference image and clip-art-like object sketches to drag and drop for scene composition. The augmentation is performed by replacing object sketches with other sketch instances belonging to the same object category. This is a dataset with sketches with a large domain gap from the freehand scene sketches we target. Yet, it is interesting to evaluate the generalization properties of our method. Tab. S5 shows a comparison of the zero-shot performance of our method (third line: Ours) with the two fully-supervised methods trained on semi-synthetic sketches. The Acc@P and mIoU are the metrics we use in the main paper. We additionally report results for two additional measures:
Mean Pixel Accuracy (MeanAcc): It measures the average pixel accuracy Acc@P of each category.
Frequency Weighted Intersection over Union (FWIoU): It introduces category occurrence frequency to the mIoU, by weighting per-category pixel IoU (intersection over union) by the frequency of occurrence.
Our model reaches high accuracy on these sketches, even in the presence of a large domain gap. In particular, the performance of our model on these sketches is higher than on the freehand and more challenging sketches from the FS-COCO dataset [8]. This, combined with the results in Tab. 3 in the main paper, is a strong argument towards usage of true freehand sketches with weak annotation in the form of captions over the semi-synthetic dataset of scene sketches.
Fine-tuning on semi-synthetic sketches
While our model does reach high accuracy on these sketches, it does not reach the performance of fully supervised methods trained on semi-synthetic sketches when tested on semi-synthetic sketches. Therefore, we investigate whether fine-tuning our model on semi-synthetic sketches can close the gap – while relying only on textual labels and not pixel-level annotations.
We perform two additional experiments:
1.
Training exclusively on Synthetic Sketches (Ours): We train our model on the SketchyScene synthetic sketches [73] using language supervision. Captions are constructed by concatenating scene sketch category names into one text token.
2.
Training on Both Synthetic and Freehand Sketches (Ours): We train the model on both SketchyScene synthetic sketches and FS-COCO freehand sketches.
The results are shown in Tab. S5: Ours and Ours.
We observe a performance increase for Ours on the sketches from the SketchyScene [73] dataset, reaching competitive performance with fully supervised methods [19, 59]. This highlights the generalization properties of our training pipeline for different data distributions and highlights that succinct captions can serve as a robust supervisory signal, lifting the need for extensive annotations.
| Method | ||||
| LDP [19] | ||||
| SketchSeger [59] | ||||
| Ours | ||||
| Ours | ||||
| Ours |
However, when freehand sketches are added to the training data (Ours), there is a slight decrease in performance across all metrics. This further emphasizes the existence of a domain gap between freehand sketches and semi-synthetic sketches, which again motivates the usage of freehand sketches with weak annotations.
| Method | |||
| LDP [19] | |||
| Ours | |||
| Ours | |||
| Ours |
Similar observations are made in Tab. S6 when the model is trained on the synthetic sketches (Ours) and tested on the FS-COCO freehand sketches. Even when both synthetic and freehand sketches are used for training (Ours), the model’s performance degrades compared to training solely on freehand sketches. This further emphasizes our observations regarding the domain gap between synthetic and freehand sketches.
| Method | Trained on | Supervision | Tested on | Segmentation accuracy | |||||
| Pixel labels | Captions | mIoU | Acc@P | Acc@C | MAcc | FWIoU | |||
| LDP [19] | SketchyScene | FS-COCO | 33.04 | 56.23 | 56.71 | 51.16 | 52.63 | ||
| SKY-Scene | ✓ | LDP freehand | 37.16 | 78.84 | - | 47.25 | 66.98 | ||
| TUB-Scene | SketchyScene | 74.93 | 93.46 | - | 85.84 | 88.13 | |||
| SketchSeger [59] | SketchyScene | FS-COCO | - | - | - | - | - | ||
| SKY-Scene | ✓ | LDP freehand | - | - | - | - | - | ||
| TUB-Scene | SketchyScene | 81.17 | 95.44 | - | 88.18 | 91.52 | |||
| Ours | FS-COCO | 73.48 | 85.54 | 87.02 | 82.27 | 84.09 | |||
| FS-COCO | ✓ | LDP freehand | 53.94 | 81.63 | - | 59.36 | 69.37 | ||
| SketchyScene | 60.91 | 87.99 | - | 66.59 | 76.33 | ||||
| Ours | SketchyScene | FS-COCO | 61.79 | 74.43 | 75.62 | 69.41 | 71.75 | ||
| ✓ | LDP freehand | 49.72 | 71.96 | - | 48.71 | 59.15 | |||
| SketchyScene | 71.73 | 92.87 | - | 79.54 | 85.19 | ||||
| Ours | FS-COCO | FS-COCO | 68.84 | 79.21 | 81.29 | 74.08 | 77.63 | ||
| SketchyScene | ✓ | LDP freehand | 50.13 | 76.07 | - | 55.83 | 62.97 | ||
| SketchyScene | 70.51 | 91.23 | - | 77.87 | 84.72 | ||||
Ours: trained on freehand sketches from the FS-COCO dataset [8], Ours is trained on synthetic sketches [73], Ours is trained on both freehand [8] and synthetic sketches [73].
We test all methods on three datasets: our FS-COCO-based test set, LDP [19] freehand sketches test set, and SketchyScene [73] synthetic sketches test set.
Training datasets: The SketchyScene [73] dataset contains 7,265 synthetic scene sketches spanning 46 categories with 5,617 images for training, and 1,113 for test. SKY-Scene and TUB-Scene were introduced in [19], and are composed of object sketches from the Sketchy [46] and TU-Berlin [14] datasets, respectively. They both have 7,265 synthetic scene sketches and follow the same data split.
Tab. S7 shows a full comparison of our method against fully supervised sketch segmentation methods: LDP [19] and SketchSeger [59], across the free datasets: FS-COCO-based test set, LDP [19] freehand sketches test set, and SketchyScene [73] synthetic sketches test set. It shows the superiority of our method on both datasets of freehand scene sketches.
S3 Detailed human study analysis
In this section, we provide a more in-depth discussion of Sec. 5: Human-model alignment of the main paper.
S3.1 Human study categories
In the main paper, in Sec. 5.1, we introduced four challenging categories of sketches for our method that we used for the user study. We show all the sketches used in the user study in Fig. S12. For convenience, below we repeat the definition of each category:
Ambiguous sketches: sketches where it might be hard even for a human observer to understand an input sketch. We selected the sketches by visually examining the test set sketches alongside reference images.
Interchangeable categories: sketches containing multiple objects with labels that can interchange each other, such as ‘tower/building’, ‘girl/man’, and ‘ground/grass’.
Correlated categories: sketches with categories that typically co-occur in scenes. These categories are semantically related. We selected sketches containing the most common pairs with significant co-occurrence. Specifically, ‘branch/bird’ (52%), ‘runway/airplane’ (44%), ‘railway/train’ (39%), and ‘road/car’ (29%), were chosen.
Numerous-categories: sketches with six or more object categories and a model accuracy (Acc@P) below . The sampled sketches have an average of categories per sketch .
Additionally, we included a Strong performance category, comprising ten sketches where the model’s accuracy (Acc@P) exceeded the average performance (85.54%), to demonstrate scenarios of effective model segmentation.
S3.2 Visual analysis of the segmentation results for the sketches with “Interchangeable categories” obtained with our model and from the user study
We conducted a visual analysis to compare the confidence in segmenting semantically similar objects by human annotators and our model. For each object category, we obtain a category confidence map by counting how many participants assigned a given label to a category. For our model, we obtain segmentation confidence as a result of a cosine similarity computation between the sketch patch features and the category textual embedding. We visualized in Fig. S13 the obtained confidence maps for the most frequently confused by our model categories: ‘girl/man’ and ‘building/tower’. We also show the pixels that are confidently assigned to belong to both considered categories (with a confidence threshold higher than ). We can observe that our model is less confident than humans in assigning labels to these categories.
S4 Additional ablation studies
S4.1 Detailed ablation on cross attention vs. self attention
To validate the effectiveness of our cross-attention module, we added a residual connection to demonstrate that relying solely on self-attention features, without the integration of cross-attention, leads to suboptimal segmentation results. We run several experiments with varying dropout ratios in the cross-attention block. This allows us to assess its impact on model performance. The results, presented in Tab. S8, show model accuracy across different dropout levels, from 0 (no dropout) to 1 (complete dropout). This shows the benefit of the design used in the main paper, equivalent to using only cross-attention in the category-level encoder.
| Dropout | 0 | 0.2 | 0.4 | 0.6 | 0.8 | 1 |
| Acc@P | 85.54 | 84.61 | 84.38 | 84.16 | 83.06 | 82.86 |
| Test set | Validation set | ||||||
| Model | Checkpoint | mIoU | Acc@P | Acc@S | mIoU | Acc@P | Acc@S |
| Optimal | 22.86 | 33.41 | 32.64 | 25.76 | 36.34 | 35.17 | |
| CLIP* | Last | 19.34 | 28.89 | 27.64 | 22.11 | 31.49 | 31.07 |
| Optimal | 45.71 | 66.21 | 66.89 | 47.26 | 68.28 | 68.76 | |
| GroupViT* | Last | 43.83 | 64.03 | 64.48 | 46.58 | 67.70 | 68.13 |
| Optimal | 49.26 | 69.87 | 73.64 | 51.27 | 71.79 | 75.67 | |
| SegCLIP* | Last | 46.41 | 66.91 | 70.31 | 50.86 | 70.12 | 74.41 |
| Optimal | 48.74 | 65.38 | 68.78 | 50.84 | 67.32 | 70.88 | |
| CLIP_Surgery* | Last | 47.29 | 63.94 | 67.13 | 48.33 | 66.01 | 68.82 |
| Optimal | 59.98 | 78.68 | 81.11 | 62.41 | 80.69 | 83.23 | |
| CLIP_Surgery** | Last | 58.64 | 77.34 | 79.88 | 61.53 | 79.41 | 82.07 |
| Optimal | 73.48 | 85.57 | 87.02 | 74.76 | 86.83 | 88.41 | |
| Ours | Last | 72.51 | 84.74 | 86.39 | 74.12 | 85.97 | 87.76 |
S4.2 Models checkpoint choice
As described in Sec. 4.3 of the main paper, for each of the models fine-tuned on sketch data: ours and competing methods, we select a checkpoint based on the performance on the validation set with pixel-lelve segmentation annotations, consisting of 475 sketches. This requires at training time having a small set of pixel-level annotated sketches, which can be limiting. However, we observe that the loss gradually decreases for our model, and it is safe to choose a last checkpoint if such an annotated set is not available. In Tab. S9, we provide a comparison with the results when for our model and competing models the last checkpoint is used. We trained for 20 epochs. We observe that after that the convergence rate is very low for each of the considered models.
We observe only a marginal performance drop (less than one point in all metrics) for our model when the last checkpoint is used compared to a checkpoint selected based on the performance on the validation set (referred to as optimal in the table). This implies that competitive model performance can be achieved without using any pixel-level annotations.
We also observe that with either of the choices of a checkpoint, the performance on the validation and test sets is similar, with just a small decrease in performance on the test set compared to the validation set. Our test set includes sketches from five non-expert artists whose sketches were not present in either the training or validation sets. Therefore, this analysis implies that there is no over-fitting to the training data and our model robustly generalizes to the unseen skecthes and drawing styles.
S4.3 Segmenting out individual categories
To explore the model’s ability to isolate individual sketch categories through thresholding, as described in Sec. 3.4 in the main paper, we assess two model versions, where (1) the optimal checkpoint is used, selected based on the performance on the validation set and (2) the last checkpoint is used (from the 20th epoch). We measure pixel accuracy (Acc@P) of segmenting a sketch into an individual category and the rest (background), employing varying threshold values. Fig. S14 shows the plot of segmentation accuracy with different threshold values on test and validation sets when either optimal Fig. S14(a.) or last Fig. S14(b.) checkpoints are used.
When optimal checkpoint is used
When using the optimal checkpoint, the model consistently achieves strong performance on validation and test sets, achieving and Acc@P, respectively, albeit at different threshold values ( and , respectively). This implies that the label assignment confidence is slightly lower on the unseen sketches in new styles. However, despite this, the model maintains a consistently strong performance on these new sketches and styles.
When the last checkpoint is used
When we use the model from the last checkpoint, the best performance on the validation and test sets is obtained with slightly lower threshold values of and , respectively. This implies that there is a correlation between the model’s confidence and its performance.

### v2

@@
 Abstract
-We study the underexplored but fundamental vision problem of machine understanding of abstract freehand scene sketches.
-We introduce a sketch encoder that results in semantically-aware feature space, which we evaluate by testing its performance on a semantic sketch segmentation task.
-To train our model we rely only on the availability of bitmap sketches with their brief captions and do not require any pixel-level annotations.
-To obtain generalization to a large set of sketches and categories, we build on a vision transformer encoder pretrained with the CLIP model.
+We study the underexplored but fundamental problem of machine understanding of abstract freehand scene sketches.
+We introduce a sketch encoder that ensures a semantically-aware feature space, which we evaluate by testing its performance on a semantic sketch segmentation task.
+To train our model, we rely only on bitmap sketches accompanied by brief captions, avoiding the need for pixel-level annotations.
+To generalize to a large set of sketches and categories, we build upon a vision transformer encoder pretrained with the CLIP model.
 We freeze the text encoder and perform visual-prompt tuning of the visual encoder branch while introducing a set of critical modifications.
-Firstly, we augment the classical key-query (k-q) self-attention blocks with value-value (v-v) self-attention blocks.
-Central to our model is a two-level hierarchical network design that enables efficient semantic disentanglement: The first level ensures holistic scene sketch encoding, and the second level focuses on individual categories.
-We, then, in the second level of the hierarchy, introduce a cross-attention between textual and visual branches.
-Our method outperforms zero-shot CLIP pixel accuracy of segmentation results by 37 points, reaching an accuracy of on the FS-COCO sketch dataset.
-Finally, we conduct a user study that allows us to identify further improvements needed over our method to reconcile machine and human understanding of scene sketches.
+First, we augment the classical key-query (k-q) self-attention blocks with value-value (v-v) self-attention blocks.
+Central to our model is a two-level hierarchical training that enables efficient semantic disentanglement: The first level ensures holistic scene sketch encoding, and the second level focuses on individual categories.
+In the second level of the hierarchy, we introduce cross-attention between the text and vision branches.
+Our method outperforms zero-shot CLIP segmentation results by 37 points, reaching a pixel accuracy of on the FS-COCO sketch dataset.
+Finally, we conduct a user study that allows us to identify further improvements needed over our method to reconcile machine and human understanding of freehand scene sketches.
 1 Introduction
@@
 1).
-We propose a two-level hierarchy in our network, where the two levels (“Holistic” and “Category-level”) share the weights of our visual encoder.
+We propose a two-level hierarchical training of our network, where the two levels (“Holistic" and “Category-level") share the weights of our visual encoder.
 The first level focuses on ensuring that our model can capture holistic scene understanding (Fig.
@@
 We also compare with a previous fully supervised work on scene sketch segmentation [21], trained on a semi-synthetic set of sketches composed of individual category sketches.
-We demonstrate that their work does not generalize well to freehand scene sketches from the FS-COCO dataset.
+We demonstrate that their work does not generalize well to freehand scene sketches [10].
 Our method demonstrates consistent performance and similarly outperforms [21] on a dataset of freehand sketches provided by Ge et al.
@@
 We then carefully assess our approach by comparing its performance with that of human participants, drawing insights to guide future work.
-In summary, our contributions include: (1) a two-level network design used during training, focusing on holistic scene sketch understanding and category disentanglement, (2) the first language-supervised scene sketch segmentation method, (3) per pixel segmentation annotations of 975 sketches from the FS-COCO dataset, and (4) multi-user annotations of a subset of distinct groups of sketches.
+In summary, our contributions include: (1) a two-level hierarchical training approach, focusing on holistic scene sketch understanding and category disentanglement, (2) the first language-supervised scene sketch segmentation method, (3) per pixel segmentation annotations of 975 sketches from the FS-COCO dataset, and (4) multi-user annotations of a subset of distinct groups of sketches.
 2 Related Work
 2.1 Unsupervised and Weekly Supervised Image Semantic Segmentation
-The need for pixel-wise segmentation limits the amount of data that fully supervised segmentation models [35, 64, 5, 6, 1, 17] can use for training, as such annotations are costly to collect.
-This in its turn limits the generalization properties of models trained with pixel-level annotations.
-To avoid the need for extensive annotations, unsupervised [62, 7, 23, 25, 38], semi-supervised [39, 72] and weakly supervised [40, 56, 58, 37, 13, 12, 36, 69] methods were proposed.
+The need for pixel-wise segmentation limits the number of instances that supervised segmentation models [38, 67, 7, 8, 1, 19] can use for training, as such annotations are costly to collect.
+This in turn limits the generalization properties of models trained with pixel-level annotations.
+To avoid the need for extensive annotations, unsupervised [65, 9, 25, 28, 41], semi-supervised [42, 75] and weakly supervised [43, 59, 61, 40, 15, 14, 39, 72, 26] methods were proposed.
 Our method belongs to the group of weakly supervised methods based on text annotations only [61, 40, 15, 14, 39, 6], such methods are not limited to a fixed set of categories and therefore are referred to as open vocabulary semantic segmentation methods.
@@
 For example, recent GroupViT [61] and SegCLIP [40] use learnable group tokens and semantic group modules to aggregate low-layer pixel features.
-In our work, we propose a two-level training architecture taking sketch specifics into account.
+In our work, we propose a two-level training architecture taking sketch sparsity and abstraction into account.
 2.2 Sketch Semantic Segmentation
@@
 All these works are fully supervised except for [44], which segments sketches of a given category provided at least one segmented reference sketch.
-Semantic scene sketch segmentation [51], and more broadly scene sketch understanding, is underexplored to a large extent due to lack of data.
+Semantic scene sketch segmentation [54], and more broadly scene sketch understanding, is underexplored, to a large extent due to a lack of data.
 The lack of data is typically addressed by introducing semi-synthetic sketch datasets.
-The SketchyScene dataset [73] consists of 7,264 sketch-image pairs, obtained by arranging clip-art sketches in alignment with a reference image.
+The SketchyScene dataset [76] consists of 7,264 sketch-image pairs, obtained by arranging clip-art individual category sketches in alignment with a reference image.
 SketchyCOCO dataset [20] is generated from COCO-Stuff [4] by semi-automatically arranging freehand sketches of individual categories.
@@
 Therefore, the input to the vision encoder is , where .
-Attention computation in the visual encoder
+Attention computation
 It was observed by Li et al.
 [35] that CLIP-predicted similarity maps between image and text features emphasize background regions rather than areas that correspond to a category in the text embedding.
-To address this issue, they proposed an alternative configuration of the visual transformer that does not require training or fine-tuning the original model.
-They demonstrated that this configuration leads to improved performance in open vocabulary segmentation tasks.
-We performed a similar experiment with CLIP features for sketch inputs: The similarity maps in the second row of Fig. 3 show the poor ability of CLIP features to identify target categories.
-Therefore, we follow Li et al.
-[32] and use their two-path configuration of the vision transformer.
-However, we use it not only for inference, but also incorporate this two-path configuration directly into our network training, as we find it more beneficial.
+To address this issue, they proposed to use an instance of self-self attention called v-v attention, which does not require training or fine-tuning the original model.
+Li et al.
+[35], and later Bousselham et al.
+[3] demonstrated that this leads to improved performance in open vocabulary segmentation tasks: Self-self-attention reinforces the similarity of tokens already close to each other (e.g.
+representing the same object), which leads to a clearer separation in the feature space, thereby improving the segmentation quality.
+We performed a similar experiment with CLIP features for sketch inputs: The similarity maps in the second row of Fig.
+3 show the poor ability of CLIP features to identify target categories.
+Therefore, we follow [35] and use their two-path configuration of the vision transformer.
+However, we use it not only for inference but also incorporate this two-path configuration directly into our network training, as we find it more beneficial.
 We provide a detailed analysis in Sec.
@@
 3.2 Categories Disentanglement
-Given the sketch caption we automatically identify individual categories and generate a set of textual prompts of the form “A sketch of *” (Fig.
+Given the sketch caption we automatically identify individual categories and generate a set of textual prompts of the form “A sketch of *" (Fig.
 2b.).
@@
 where is the patch index and is an index of a category (e.g.
-tree).
+trees).
 The resulting similarity matrix represents the category label probabilities for each individual patch (Fig.
@@
 2e., we obtain disentanglement into individual sketch categories.
-3.2.1 Thresholding with a learnable parameter
+Thresholding with a learnable parameter
 Only pixels with similarity scores above a certain threshold are retained at this step (Fig.
@@
 More importantly, the threshold value increases over epochs as the model becomes more confident in its predictions, allowing the model to obtain strong disentanglement performance.
-3.2.2 Visual Encoder with Cross Attention
+Visual encoder with Cross-Attention
 The features of individual category sketches are extracted with the visual encoder identical to the one used in the holistic scene sketch level understanding of our network, described in Sec.
@@
 4 underscores the efficacy of this cross-attention strategy.
-3.2.3 Text-sketch category-level alignment
+Text-sketch category-level alignment
 We train with a triplet loss, , so that the category sketch embedding, VCT (Vision Category Token), is used as an anchor, the matching embedding of the category prompt is used as a positive sample and the embedding of the prompt of the most dissimilar category is used as negative.
@@
 We do not use the learned value from the training, as during training the model does not have to select all the pixels of the given category, but only those that are sufficient to confidently predict the category label.
-We provide in-depth discussion in the supplemental.
+We provide an in-depth discussion in the supplemental.
 4 Experiments
@@
 We use the remaining 9025 sketches for training.
-Test/Validation set annotations One of the co-authors manually annotated test and validation sketches, relying on reference images and category labels from the MS-COCO [33] dataset.
+Annotations One of the co-authors manually annotated test and validation sketches, relying on reference images and category labels from the MS-COCO [36] dataset.
 We assign each stroke a unique category label.
@@
 4.4 Comparison against state-of-the-art
-4.4.1 Comparison with methods relying on the availability of pixel-level annotations
+4.4.1 Comparison with fully-supervised methods
 We first compare with several recent methods for image segmentation that similarly to us utilize either CLIP as a backbone: DenseCLIP [47] and ZegCLIP [72], or more recent foundational backbones Grounding-DINO [37] and SAM [31], used in Grounded-SAM [22].
-These methods require a certain amount of pixel-level annotated examples.
-We also compare to a recent fully supervised method [19] for scene sketch semantic segmentation, which is trained on a dataset of semi-synthetic sketches.
+These methods require pixel-level annotated examples, and therefore can not be fine-tuned on our training data.
+We also compare to a recent fully supervised method LDP (Local Detail Perception) [21] for scene sketch semantic segmentation, which is trained on a dataset of semi-synthetic sketches.
 Such sketches are obtained as a superposition of freehand category-level sketches.
-We refer to this method as LDP (Local Detail Perception).
-Tab. 1 shows that neither of the considered methods generalizes well to freehand scene sketches with a large number of categories.
+Tab.
+1 shows that neither of the these methods generalizes well to freehand scene sketches.
 | Methods | |||
@@
 The fine-tuned methods are marked with stars.
-Table 2 shows that our method outperforms all considered baselines, and surpasses the best-performing baseline CLIP Surgery by a substantial margin of , and points in score, and , respectively.
+Tab.
+2 shows that our method outperforms all considered baselines, and surpasses the best-performing baseline CLIP Surgery by a substantial margin of , and points in score, and , respectively.
 In Sec.
 4.5.1, we evaluate various elements of our architecture and their contribution to overall performance.
-| Methods | |||
-| CLIP [43] | |||
-| GroupViT [58] | |||
-| SegCLIP [37] | |||
-| CLIP_Surgery [32] | |||
-| CLIP | |||
-| GroupViT | |||
-| SegCLIP | |||
-| CLIP_Surgery | |||
-| CLIP_Surgery | |||
+| Methods | ||||
+| Zero-shot | CLIP [46] | |||
+| GroupViT [61] | ||||
+| SegCLIP [40] | ||||
+| CLIP_Surgery [35] | ||||
+| Fine-tuned | CLIP | |||
+| GroupViT | ||||
+| SegCLIP | ||||
+| CLIP_Surgery | ||||
+| CLIP_Surgery | ||||
 | Ours | 73.48 | 85.54 | 87.02 |
@@
 4.4.3 Generalization ability of our method
-Here, we evaluate our method on an additional dataset of 50 freehand sketches provided and annotated by Ge et al.
+Next, we evaluate our method on an additional dataset of 50 freehand sketches provided and annotated by Ge et al.
 [21].
@@
 The lower mIoU values on these sketches than on FS-COCO sketches can be explained by (1) on larger average number of categories in them ( categories per sketch) than in our FS-COCO test set ( categories per sketch); (2) domain gap.
-The sketches from [19] contain symbolic representations of objects (see inset
+The sketches from [21] contain symbolic representations of objects (see the inset
 4.5 Ablation Study
@@
 We sample 5 sketches for each of the first 4 categories and 10 sketches for the 5th category. We visualize selected sketches in the supplemental material.
-Participants We recruited 25 participants ( male) for this study.
-Each participant was randomly assigned six sketches: 1 from each of the first 4 groups and 2 from the 5th group, such that every sketch was annotated by five unique participants.
+Participants We recruited 25 participants ( male).
+Each participant was randomly assigned 6 sketches: 1 from each of the first 4 groups and 2 from the 5th group, such that every sketch was annotated by five unique participants.
 Study Procedure Participants were presented with one sketch and one object category at a time and were not able to see their previous annotations.
@@
 8 Strong).
-Interestingly, given a label humans can correctly identify sketch pixels even in the presence of ambiguity (Fig.
-8 Ambiguous).
+Fig.
+8 Ambiguous shows that, given a label, humans can correctly identify sketch pixels even in the presence of ambiguity.
 While none of the models currently match human performance on ambiguous sketches, our model surpasses the other methods by a noticeable margin, demonstrating the effectiveness of our two-level training architecture.
@@
 In Sec. S4.3, we discuss in detail the choice of a threshold value for segmenting out pixels corresponding to individual categories.
-S2 Additional analysis of our method performance
+In Sec. S5, we provide the computational cost of our method.
+S2 Additional Performance Analysis
 S2.1 Additional Qualitative Comparisons
@@
 S11 the pixel accuracy (Acc@P) for each selected object category.
-For the figure, we selected categories that appear more than ten times in FS-COCO dataset [8] captions.
+For the figure, we selected categories that appear more than ten times in the FS-COCO dataset [10] captions.
 First, we can see that the segmentation accuracy is smoothly distributed across different categories.
@@
 To evaluate this, we approximate the frequency of a category by counting its occurrence in both the train and test sets, then consider only categories that appear in the test set.
-We plot with a green and red lines in Fig.
+We plot with green and red lines in Fig.
 S11 the train and test sets category frequency, respectively.
@@
 Therefore, the model efficiently leverages pre-training on a large image dataset.
-S2.3 Additional discussion on synthetic vs. freehand sketches
+Model generalization to new object categories
+Our test set includes 185 object classes, with 125 seen and 60 unseen during training.
+The accuracy on seen categories is and on unseen.
+These results demonstrate good generalization of our model to unseen categories.
+S2.3 Synthetic vs. Freehand sketches
 In the Sec.
@@
 3 in the main paper, is a strong argument towards usage of true freehand sketches with weak annotation in the form of captions over the semi-synthetic dataset of scene sketches.
-Fine-tuning on semi-synthetic sketches
-While our model does reach high accuracy on these sketches, it does not reach the performance of fully supervised methods trained on semi-synthetic sketches when tested on semi-synthetic sketches.
-Therefore, we investigate whether fine-tuning our model on semi-synthetic sketches can close the gap – while relying only on textual labels and not pixel-level annotations.
-We perform two additional experiments:
-1.
-Training exclusively on Synthetic Sketches (Ours): We train our model on the SketchyScene synthetic sketches [73] using language supervision.
-Captions are constructed by concatenating scene sketch category names into one text token.
-2.
-Training on Both Synthetic and Freehand Sketches (Ours): We train the model on both SketchyScene synthetic sketches and FS-COCO freehand sketches.
-The results are shown in Tab. S5: Ours and Ours.
-We observe a performance increase for Ours on the sketches from the SketchyScene [73] dataset, reaching competitive performance with fully supervised methods [19, 59].
-This highlights the generalization properties of our training pipeline for different data distributions and highlights that succinct captions can serve as a robust supervisory signal, lifting the need for extensive annotations.
 | Method | ||||
@@
 | Ours |
-However, when freehand sketches are added to the training data (Ours), there is a slight decrease in performance across all metrics.
-This further emphasizes the existence of a domain gap between freehand sketches and semi-synthetic sketches, which again motivates the usage of freehand sketches with weak annotations.
-| Method | |||
-| LDP [19] | |||
-| Ours | |||
-| Ours | |||
-| Ours |
-Similar observations are made in Tab.
-S6 when the model is trained on the synthetic sketches (Ours) and tested on the FS-COCO freehand sketches.
-Even when both synthetic and freehand sketches are used for training (Ours), the model’s performance degrades compared to training solely on freehand sketches.
-This further emphasizes our observations regarding the domain gap between synthetic and freehand sketches.
 | Method | Trained on | Supervision | Tested on | Segmentation accuracy | |||||
@@
 They both have 7,265 synthetic scene sketches and follow the same data split.
+| Method | |||
+| LDP [21] | |||
+| Ours | |||
+| Ours | |||
+| Ours |
+Fine-tuning on semi-synthetic sketches
+While our model does reach high accuracy on these sketches, it does not reach the performance of fully supervised methods trained on semi-synthetic sketches when tested on semi-synthetic sketches.
+Therefore, we investigate whether fine-tuning our model on semi-synthetic sketches can close the gap – while relying only on textual labels and not pixel-level annotations.
+We perform two additional experiments:
+1.
+Training exclusively on Synthetic Sketches (Ours): We train our model on the SketchyScene synthetic sketches [76] using language supervision.
+Captions are constructed by concatenating scene sketch category names into one text token.
+2.
+Training on Both Synthetic and Freehand Sketches (Ours): We train the model on both SketchyScene synthetic sketches and FS-COCO freehand sketches.
+The results are shown in Tab. S5: Ours and Ours.
+We observe a performance increase for Ours on the sketches from the SketchyScene [76] dataset, reaching competitive performance with fully supervised methods [21, 62].
+This highlights the generalization properties of our training pipeline for different data distributions and highlights that succinct captions can serve as a robust supervisory signal, lifting the need for extensive annotations.
+However, when freehand sketches are added to the training data (Ours), there is a slight decrease in performance across all metrics.
+This further emphasizes the existence of a domain gap between freehand sketches and semi-synthetic sketches, which again motivates the usage of freehand sketches with weak annotations.
+Similar observations are made in Tab.
+S7 when the model is trained on the synthetic sketches (Ours) and tested on the FS-COCO freehand sketches.
+Even when both synthetic and freehand sketches are used for training (Ours), the model’s performance degrades compared to training solely on freehand sketches.
+This further emphasizes our observations regarding the domain gap between synthetic and freehand sketches.
 Tab.
-S7 shows a full comparison of our method against fully supervised sketch segmentation methods: LDP [19] and SketchSeger [59], across the free datasets: FS-COCO-based test set, LDP [19] freehand sketches test set, and SketchyScene [73] synthetic sketches test set.
+S6 shows a full comparison of our method against fully supervised sketch segmentation methods: LDP [21] and SketchSeger [62], across the free datasets: FS-COCO-based test set, LDP [21] freehand sketches test set, and SketchyScene [76] synthetic sketches test set.
 It shows the superiority of our method on both datasets of freehand scene sketches.
+S2.3.2 Pre-training on synthetic sketches
+We also experiment with fine-tuning CLIP and CLIPSurgery on synthetic sketches.
+However, training on millions of synthetic sketches is out of the scope of this work due to computational constraints.
+As a feasible experiment, we generated synthetic sketches for the reference images in our training set, using [5], in ‘contour’ style (as the closest to the test set sketches style).
+This is the number of sketches identical to the number of sketches we use to train our model.
+The accuracy on our test set of fine-tuned this way CLIP and CLIPSurgery increases by negligible to points compared to their zero-shot performance.
+In comparison, our model outperforms their zero-shot performance by and points, respectively.
+Training our model from CLIP weights pre-trained on synthetic sketches boosts the performance only by points.
 S3 Detailed human Study Analysis
@@
 In the main paper, in Sec.
-5.1, we introduced four challenging categories of sketches for our method that we used for the user study.
+5.1, we introduced four challenging categories of sketches for our method, that we used for the user study.
 We show all the sketches used in the user study in Fig.
@@
 Additionally, we included a Strong performance category, comprising ten sketches where the model’s accuracy (Acc@P) exceeded the average performance (85.54%), to demonstrate scenarios of effective model segmentation.
-S3.2 Visual analysis of the segmentation results for the sketches with “Interchangeable categories” obtained with our model and from the user study
+S3.2 Annotators
+We recruited 25 participants ( male).
+The annotators are PhD students in diverse disciplines and of diverse nationalities aged from 22 to 42 years (average age 29.32).
+We believe this group represents well the general population and each individual performed the task carefully.
+S3.3 Visual Analysis of Interchangeable Categories Segmentation Results
 We conducted a visual analysis to compare the confidence in segmenting semantically similar objects by human annotators and our model.
@@
 We can observe that our model is less confident than humans in assigning labels to these categories.
+S3.4 Statistical Significance: Ours vs CLIPSurgery
+On the 20 sketches from the 4 challenging groups, our model outperforms CLIPSurgery with a p-value of .
+In the ‘strong’ group, we have 10 sketches, in which, while our model performs on par with humans, it outperforms CLIPSurgery with a p-value of .
 S4 Additional Ablation Studies
@@
 As described in Sec.
-4.3 of the main paper, for each of the models fine-tuned on sketch data: ours and competing methods, we select a checkpoint based on the performance on the validation set with pixel-lelve segmentation annotations, consisting of 475 sketches.
+4.3 of the main paper, for each of the models fine-tuned on sketch data: ours and competing methods, we select a checkpoint based on the performance on the validation set with pixel-level segmentation annotations, consisting of 475 sketches.
 This requires at training time having a small set of pixel-level annotated sketches, which can be limiting.
@@
 Our test set includes sketches from five non-expert artists whose sketches were not present in either the training or validation sets.
-Therefore, this analysis implies that there is no over-fitting to the training data and our model robustly generalizes to the unseen skecthes and drawing styles.
+Therefore, this analysis implies that there is no over-fitting to the training data and our model robustly generalizes to the unseen sketches and drawing styles.
+| Training | Parameters | Inference | |||||
+| GFLOPS | Hours | Epochs | # Trainable | # Full | GFLOPS | Secs. | |
+| CLIP | 189.47 | 3.76 | 20 | 149M | 149M | 89.02 | 0.21 |
+| CLIPSurgery | 231.41 | 4.08 | 20 | 162M | 162M | 113.64 | 0.73 |
+| Ours | 346.36 | 7.31 | 20 | 16M | 165M | 121.59 | 1.05 |
 S4.3 Segmenting out Individual Categories
@@
 This implies that there is a correlation between the model’s confidence and its performance.
+S5 Computational Cost
+We detail in Tab.
+S10 the computational cost of our method compared to CLIP and CLIPSurgery.
+Our two-level hierarchical network design introduces additional complexity, through value-value self-attention and cross-attention blocks.
+However, we maintain a comparable level of complexity to CLIPSurgery during inference.
+This slight computational increase is justified given the substantial 13 mIoU points improvement over CLIP_Surgery (as shown in Tab.
+4 in the main paper).
+Our code can be further optimized to reach the performance of CLIPSurgery at inference time.